The most recent financial statements for Cardinal, Inc., are shown here:

Income Statement Balance Sheet
  Sales $ 33,000   Assets $ 77,200   Debt $ 40,200
  Costs 18,650   Equity 37,000
  Taxable income $ 14,350     Total $ 77,200     Total $ 77,200
  Taxes (24%) 3,444
    Net income $ 10,906

Assets and costs are proportional to sales. Debt and equity are not. A dividend of $4,200 was paid, and the company wishes to maintain a constant payout ratio. Next year’s sales are projected to be $37,950.

What is the external financing needed? (Do not round intermediate calculations.)
Answer #1

Dividend payout ratio=Dividend/Net income

=(4200/10906)=0.385109114

Growth rate in sales=(37950-33000)/33000=15%

Sales 37950
Costs(18650*1.15) 21447.5
Taxable income 16502.5
Taxes(16502.5*24%) 3960.6
Net income $12541.9
Less:Dividends($12541.9*0.385109114) 4830
Addition to retained earnings $7711.9

Total assets would be=$77200*1.15=$88780

Total equity would be=$37000+Addition to retained earnings

=(37000+7711.9)=$44711.9

Total assets=Total equity+Total liabilities

Hence external financing needed=$88780-(44711.9+40200)

=$3868.1